# Causal structure and topology change in (2+1)-dimensional simplicial gravity

### v1

@@ Abstract
We develop a systematic method for analyzing the causal structure at vertices in (2+1)–dimensional Lorentzian simplicial gravity. By examining the intersection patterns of lightcones emanating from a vertex with its simplicial neighbourhood, we identify 13 distinct causal types of Lorentzian tetrahedra—excluding configurations with null faces. This forms the basis for a topological classification of the local causal structure in terms of the number of connected spacelike and timelike regions on the triangulated 2-sphere surrounding the vertex. These local causal data allow us to identify regular causal structures and new types of causal irregular configurations, generalizing well-known (1+1)–dimensional topologies, such as ‘Trousers’ and ‘Yarmulkes’, to higher dimensions.
We further investigate the dynamical implications of vertex causality by analyzing the behaviour of deficit angles and the Regge action in explicit configurations. Transitions in vertex causal structure coincide with singularities in the curvature and action, suggesting discrete topology change. Causally irregular hinges correspond to discrete conical singularities, while vertex causal irregularities manifest as point-like singularities. Interestingly, we find that vertex and hinge causality are generally independent. These results have direct implications for discrete quantum gravity approaches where the emergence of semiclassical spacetime depends on how causal structures are encoded within the triangulation.
@@ 1 Introduction
In Lorentzian Regge calculus, the metric structure is replaced by edge lengths assigned to simplices, and the curvature is encoded in deficit angles concentrated at the hinges (sub-simplices of codimension 2), much like in the Euclidean case. These Lorentzian deficit angles do not only encode curvature but also determine the causal structure at the hinges. Explicitly, they count how many lightrays cross the regions around the two dimensional subspace orthogonal to a given hinge [11, 12, 13]. These configurations may be viewed as discrete analogues of ‘conical singularities’ in the continuum [14, 15], where lightcone structure is disrupted along codimension-2 defects. While the role of hinges has been extensively studied [16, 12, 17, 18, 19], a similar characterization of the causal structure at lower-dimensional subsimplices (such as edges and vertices in higher dimensions) has remained largely unexplored. In this article, we address this gap by providing a detailed analysis of causal structure at vertices in (2+1)–dimensional Lorentzian triangulations. Causally irregular vertices can be interpreted as ‘point-like singularities’ (discrete analogues of continuum loci where the Lorentzian metric becomes ill-defined), often corresponding to topological transitions.
@@
In this article, we develop an intrinsic, coordinate–free approach that captures the local causal structure of a bulk vertex purely from its surrounding simplicial geometry in 2+1 dimensions. Similar notions have appeared in [34, 35] for a restricted class of simplicial geometries used in CDT. Our central idea to determine the vertex causal structure is to analyze how the lightcones at a vertex intersect its surrounding simplicial neighbourhood. We introduce a classification of Lorentzian tetrahedra according to the intersection patterns between the lightcones and the face opposite the bulk vertex. This yields 13 distinct causal types of tetrahedra when null edges or faces are excluded (see Proposition 1). Each type corresponds to a unique configuration of lightcone intersection curves, typically planar conic sections, on the opposite triangles. More generally, in 2+1 dimensions, the neighbourhood of a bulk vertex in a triangulation is homeomorphic to a three-ball, with its boundary given by a piecewise-flat 2-sphere $S^{2}$ comprised of the triangles opposite the central vertex. Analyzing the lightcone intersection patterns across this boundary sphere allows one to define a natural set of topological invariants, which enumerate the number of connected spacelike and timelike regions on the surface of the triangulated sphere. These topological invariants encode the local causal structure at the vertex and form the foundation for classifying both regular and irregular causal structures in Lorentzian triangulations.
@@
The remainder of the paper is structured as follows: In Section 2, we review the Lorentzian geometry of a simplicial complex, introducing generalized simplex inequalities and discussing the classical dynamics governed by the Regge action in $2+1$ dimensions. Section 3 analyzes the lightcone structure associated with a sub-simplex in a Lorentzian triangulation, focusing on how its local causal structure is determined by the intersection of lightcones with its simplicial neighbourhood. In Section 4, we present a classification of the causal types of Lorentzian tetrahedra based on the intersection patterns of the lightcones at a vertex with its opposite triangle. This allows us to formulate a procedure for determining the local causal structure of a bulk vertex from the counts of connected spacelike and timelike regions on its boundary sphere. Section 5 explores the dynamical implications of these causal structures via the behaviour of deficit angles and the Regge action. We conclude in Section 6 with a discussion of the broader significance of these results and potential directions for future research.
@@ 2 Lorentzian simplicial geometry
To describe the geometry of a simplex, place an arbitrary vertex, say $v_{0}$, at the origin. Then, the set of edge vectors $\{\vec{e}_{i}\}_{i=1}^{d}$ based at $v_{0}$, where $\vec{e}_{i}=v_{i}-v_{0}$, completely characterize the geometry of the simplex. From the edge vectors the squared edge lengths, angles and volumes of faces of the simplex can be determined. For instance, the squared length associated with an edge with edge vector $\vec{e}_{i}$ is given by the Minkowski inner product $s_{i}=\langle\vec{e}_{i},\vec{e}_{i}\,\rangle\equiv\eta_{ab}e^{a}_{i}e^{b}_{i}$, where $\eta={\rm diag}(-1,1,\dots,1)$. The edge lengths can therefore be spacelike, timelike or null depending on the sign of its squared length. The inner product between a pair of edge vectors $\vec{e}_{i}$ and $\vec{e}_{j}$, which encodes the angle between them, can be expressed in terms of edge lengths111 For an Euclidean simplex, the standard Euclidean inner product between the edge vectors gives the same expression in terms of the squared lengths. as
$\langle\vec{e}_{i},\vec{e}_{j}\,\rangle=\tfrac{1}{2}(s_{i}+s_{j}-s_{ij})\,,\quad\quad$ (1)
@@
A set of $\tfrac{1}{2}d(d+1)$ real numbers $\{s_{ij}\}_{1\leq i,j\leq d}$ is realizable as the set of squared edge lengths of a simplex in Minkowski spacetime $\mathbb{R}^{1,d-1}$ if and only if the corresponding Gram matrix $G(s)$ has signature (,+,,+)(-,+$\dots$,+)( - , + , … , + ), i.e. it has exactly one negative and $d-1$ positive eigenvalues.
A proof of Proposition 1 is provided in appendix A. The generalized Lorentzian simplex inequalities impose algebraic constraints on the squared edge lengths of a simplex. These constraints can also be expressed in terms of the signed volumes of the sub-simplices, computed using the Cayley–Menger determinant [60, 29]. Generally, the inequalities requires that if a sub-simplex is spacelike, then all of its faces must also be spacelike. In contrast, a timelike sub-simplex may contain a mixture of spacelike, timelike, or null faces.
@@ 2.1 Lorentzian Regge Calculus
Regge calculus [1] provides a discretization of general relativity in which curvature is localized on co-dimension two subsimplices (commonly referred to as hinges) within a piecewise-flat manifold called a triangulation. In Lorentzian signature, the building blocks are taken to be flat Minkowski simplices, and geometric quantities such as angles and volumes of simplices are adapted to the causal and algebraic structure of Minkowski spacetime. We focus here on the (2+1)–dimensional case, which captures the essential features of Lorentzian simplicial geometry relevant to our study. In 2+1 dimensions, the fundamental building blocks are flat Lorentzian tetrahedra (3-simplices) embeddedable in $\mathbb{R}^{1,2}$. Curvature is concentrated on 1-dimensional edges of the fixed triangulation $\mathcal{T}$ and encoded via deficit angles around each edge. In Lorentzian signature, the dihedral angle at an edge depends on whether the edge is timelike or spacelike [11, 12, 13]. For timelike edges, the dihedral angles between adjacent faces are given by the standard Euclidean formulas and are always real. In contrast, for spacelike edges, the dihedral angles are determined by Lorentzian geometry and can take real or complex values. A careful treatment is therefore necessary to account for these causal distinctions and to ensure that the Regge action remains well-defined.
@@
Here [^2] The expressions for $N_{a},N_{b}$ and $N_{ab}$ are also the inner products of normal vectors associated with the triangles $t_{a},t_{b}$. The dihedral angles are defined as the angle between the outward-pointing normals of two adjacent faces of the tetrahedron, and are measured in the plane orthogonal to the edge (hinge)., $\mathbb{V}_{\tau}$ is the squared volume of the tetrahedron, $\mathbb{A}_{t_{a}}$ is the squared area of the triangle $t_{a}$, $s_{e}$ is the squared length of the edge $e$, and $s_{\bar{e}}$ is the squared length of the edge opposite to the edge $e$. For the complex dihedral angle defined in Eq.(3), we adopt the convention that $\sqrt{-r}=\imath\sqrt{r}$ and $\log_{-}(-r)=\log r-\imath\pi$ for $r\geq 0$, where the branch cuts of the square root and logarithm are on the negative real axis. All quantities in Eq. (4) can be expressed purely in terms of edge lengths of the tetrahedron, e.g. via Cayley-Menger determinant for the simplices [60, 29].
The complex dihedral angles in Eq. (3) capture the definitions of both Euclidean and Lorentzian angles. It naturally generalizes planar angles to higher dimensions. For a timelike edge within a Lorentzian tetrahedron, the corresponding dihedral angle is defined via Euclidean angles given by $\theta_{e,\tau}^{E}:=-\Theta_{e,\tau}\in[0,\pi]$. For a spacelike edge, the complex dihedral angle defines the corresponding Lorentzian dihedral angle which takes the form $\theta_{e,\tau}^{L}:=\Theta_{e,\tau}=\imath\beta-m\pi/2$, where $\beta\in\mathbb{R}$ and $m\in\{0,1,2\}$. The imaginary part $\beta$ corresponds to the rapidity between the adjacent face normals, characterizing the hyperbolic angle associated with the Lorentz boost. The integer $m$, which contributes to the real part of the dihedral angle, counts the number of lightrays that pass through the convex wedge formed by the two adjacent faces. This wedge lies in the 2D plane orthogonal to the edge $e$ where the dihedral angle is defined. Geometrically, $m$ contributes to the causal structure of the edge by counting the number of lightrays intersecting its orthogonal plane.
@@ 3 Lightcone analysis in Lorentzian triangulation
A defining feature of a Lorentzian geometry is the presence of a lightcone structure, which locally divides the region around each spacetime point into causally connected and disconnected parts (see Figure 1(a)). Unlike in the continuum, where local lightcones are encoded in the metric, discrete triangulations do not explicitly encode causal structure. Instead, this causal information must be reconstructed from the simplicial geometry. This section presents a combinatorial and algebraic approach to reconstructing lightcone data at a vertex from edge lengths in a (2+1)–dimensional Lorentzian triangulation. We shall mainly focus on Lorentzian simplices with no null edges or faces, i.e., each edge or triangle of a tetrahedron is either spacelike or timelike. The lightcone analysis at vertices within a Lorentzian triangulation can easily be extended to accommodate null edges or null faces.
@@
(b) A Lorentzian tetrahedron $\Delta(v_{0}v_{1}v_{2}v_{3})$ with its reference vertex $v_{0}$ at the origin intersecting the lightcones at $v_{0}$ (in orange). The intersection traces out null curves as conic sections on the opposite triangle $\Delta(v_{1}v_{2}v_{3})$.
Figure 1: Lightcones at a vertex in (2+1)–dimensional Minkowksi spacetime.
3.1 Local causality conditions in Lorentzian triangulation

### v2

@@ Abstract
We develop a systematic method for analyzing the causal structure at vertices in (2+1)–dimensional Lorentzian simplicial gravity. By examining the intersection patterns of lightcones emanating from a vertex with its simplicial neighbourhood, we identify 13 distinct causal types of Lorentzian tetrahedra—excluding configurations with null faces. This classification forms the basis for a topological characterization of the local causal structure in terms of the number of connected regions on the triangulated 2-sphere that are spacelike and timelike separated from a bulk vertex. These local causal data allow us to identify regular causal configurations and new types of irregular causal configurations, generalizing well-known (1+1)–dimensional topologies, such as ‘Trousers’ and ‘Yarmulkes’, to higher dimensions.
We further investigate the dynamical implications of vertex causality by analyzing the behaviour of deficit angles and the Regge action in explicit configurations. Transitions in ‘vertex–causality’ coincide with discontinuities in the curvature, suggesting discrete topology change. Causally irregular hinges correspond to discrete conical singularities, while vertex causal irregularities manifest as point-like singularities. Interestingly, we find that vertex and hinge causality are generally independent. These results have direct implications for discrete quantum gravity approaches where the emergence of semiclassical spacetime depends on how causal structures are encoded within the triangulation.
@@ 1 Introduction
In Lorentzian Regge calculus, the metric structure is replaced by edge lengths assigned to simplices, and the curvature is encoded in deficit angles concentrated at the hinges (subsimplices of codimension–2), much like in the Euclidean case. These Lorentzian deficit angles do not only encode curvature but also determine the causal structure at the hinges. Explicitly, they count how many lightrays cross the regions around the two dimensional subspace orthogonal to a given hinge [11, 12, 13]. These configurations may be viewed as discrete analogues of ‘conical singularities’ in the continuum [14, 15], where lightcone structure is disrupted along codimension–2 defects. While the role of hinges has been extensively studied [16, 12, 17, 18, 19], a similar characterization of the causal structure at lower-dimensional subsimplices (such as edges and vertices in higher dimensions) has remained largely unexplored. In this article, we address this gap by providing a detailed analysis of causal structure at vertices in (2+1)–dimensional Lorentzian triangulations. Causally irregular vertices can be interpreted as ‘point-like singularities’ (discrete analogues of continuum loci where the Lorentzian metric becomes ill-defined), often corresponding to topological transitions.
@@
In this article, we develop an intrinsic, coordinate–free approach that captures the local causal structure of a bulk vertex purely from its surrounding simplicial geometry in 2+1 dimensions. Similar notions have appeared in [34, 35] for a restricted class of simplicial geometries used in CDT. Our central idea is to determine the vertex–causal structure by analyzing how the lightcones centered at a vertex intersect its surrounding simplicial neighbourhood. We introduce a classification of Lorentzian tetrahedra based on the intersection patterns between the lightcones and the face opposite the bulk vertex. This yields 13 distinct causal types of tetrahedra when null edges or faces are excluded (see Proposition 1). Each type corresponds to a unique configuration of lightcone intersection curves, typically planar conic sections, on the opposite triangles. More generally, in 2+1 dimensions, the neighbourhood of a bulk vertex in a triangulation is homeomorphic to a three-ball, with its boundary given by a piecewise-flat 2-sphere $S^{2}$ comprised of the triangles opposite the central vertex. The intersections of the lightcones centered at the vertex with the boundary sphere partition the sphere into regions that are spacelike or timelike with respect to that vertex. That is, regions whose points are spacelike or timelike separated from the vertex according to the local Minkowski geometry of each incident tetrahedron. Analyzing the connectivity of these regions defines a natural set of topological invariants, which count the number of connected spacelike and timelike domains relative to the vertex. These invariants characterize the local causal structure and provide the foundation for distinguishing regular and irregular vertex configurations in Lorentzian triangulations.
@@
The remainder of the paper is structured as follows: In Section 2, we review the Lorentzian geometry of a simplicial complex, introducing generalized simplex inequalities and discussing the classical dynamics governed by the Regge action in $2+1$ dimensions. Section 3 analyzes the lightcone structure associated with a subsimplex in a Lorentzian triangulation, focusing on how its local causal structure is determined by the intersection of lightcones with its simplicial neighbourhood. In Section 4, we present a classification of the causal types of Lorentzian tetrahedra based on the intersection patterns of the lightcones at a vertex with its opposite triangle. This allows us to formulate a procedure for determining the local causal structure of a bulk vertex by counting the number of connected regions on the boundary sphere surrounding the vertex that are spacelike and timelike with respect to the vertex. Section 5 explores the dynamical implications of these causal structures via the behaviour of deficit angles and the Regge action. We conclude in Section 6 with a discussion of the broader significance of these results and potential directions for future research.
@@ 2 Lorentzian simplicial geometry
To describe the geometry of a simplex, place an arbitrary vertex, say $v_{0}$, at the origin. Then, the set of edge vectors $\{\vec{e}_{i}\}_{i=1}^{d}$ based at $v_{0}$, where $\vec{e}_{i}=v_{i}-v_{0}$, completely characterize the geometry of the simplex. From the edge vectors the squared edge lengths, angles and volumes of faces of the simplex can be determined. For instance, the squared length associated with an edge with edge vector $\vec{e}_{i}$ is given by the Minkowski inner product $s_{i}=\langle\vec{e}_{i},\vec{e}_{i}\,\rangle\equiv\eta_{ab}e^{a}_{i}e^{b}_{i}$, where $\eta={\rm diag}(-1,1,\dots,1)$. The edge lengths can therefore be spacelike, timelike or null depending on the sign of its squared length. The inner product between a pair of edge vectors $\vec{e}_{i}$ and $\vec{e}_{j}$, which encodes the angle between them, can be expressed in terms of edge lengths [^1] For an Euclidean simplex, the standard Euclidean inner product between the edge vectors gives the same expression in terms of the squared lengths. as
$\langle\vec{e}_{i},\vec{e}_{j}\,\rangle=\tfrac{1}{2}(s_{i}+s_{j}-s_{ij})\,$$\quad\quad$ (1)
@@
A set of $\tfrac{1}{2}d(d+1)$ real numbers $\{s_{ij}\}_{1\leq i,j\leq d}$ is realizable as the set of squared edge lengths of a simplex in Minkowski spacetime $\mathbb{R}^{1,d-1}$ if and only if the corresponding Gram matrix $G(s)$ has signature (,+,,+)(-$\dots,+$), i.e. it has exactly one negative and d1d-1 positive eigenvalues.
A proof of Proposition 1 is provided in Appendix A. The generalized Lorentzian simplex inequalities impose algebraic constraints on the squared edge lengths of a simplex. These constraints can also be expressed in terms of the signed volumes of the subsimplices, computed using the Cayley–Menger determinant [60, 29]. Generally, the inequalities requires that if a subsimplex is spacelike, then all of its faces must also be spacelike. In contrast, a timelike subsimplex may contain a mixture of spacelike, timelike, or null faces.
@@ 2.1 Lorentzian Regge Calculus
Regge calculus [1] provides a discretization of general relativity in which curvature is localized on codimension–2 subsimplices (commonly referred to as hinges) within a piecewise-flat manifold called a triangulation. In Lorentzian signature, the building blocks are taken to be flat Minkowski simplices, and geometric quantities such as angles and volumes of simplices are adapted to the causal and algebraic structure of Minkowski spacetime. We focus here on the (2+1)–dimensional case, which captures the essential features of Lorentzian simplicial geometry relevant to our study. In 2+1 dimensions, the fundamental building blocks are flat Lorentzian tetrahedra (3-simplices) embeddable in $\mathbb{R}^{1,2}$. Curvature is concentrated on 1-dimensional edges of the fixed triangulation $\mathcal{T}$ and encoded via deficit angles around each edge. In Lorentzian signature, the dihedral angle at an edge depends on whether the edge is timelike or spacelike [11, 12, 13]. For timelike edges, the dihedral angles between adjacent faces are given by the standard Euclidean formulas and are always real. In contrast, for spacelike edges, the dihedral angles are determined by Lorentzian geometry and can take real or complex values. A careful treatment is therefore necessary to account for these causal distinctions and to ensure that the Regge action remains well-defined.
@@
Here [^2] The expressions for $N_{a},N_{b}$ and $N_{ab}$ are also the inner products of normal vectors associated with the triangles $t_{a},t_{b}$. The dihedral angles are defined as the angle between the outward-pointing normals of two adjacent faces of the tetrahedron, and are measured in the plane orthogonal to the edge (hinge)., $\mathbb{V}_{\tau}$ is the squared volume of the tetrahedron, $\mathbb{A}_{t_{a}}$ is the squared area of the triangle $t_{a}$, $s_{e}$ is the squared length of the edge $e$, and $s_{\bar{e}}$ is the squared length of the edge opposite to the edge $e$. For the complex dihedral angle defined in Eq.(3), we adopt the convention that $\sqrt{-r}=\imath\sqrt{r}$ and $\log_{-}(-r)=\log r-\imath\pi$ for $r\geq 0$, where the branch cuts of the square root and logarithm are on the negative real axis. All quantities in Eq. (4) can be expressed purely in terms of edge lengths of the tetrahedron. For a triangle $t$ with squared edge lengths $\{s_{a},s_{b},s_{c}\}$, the squared area is simply given by Heron’s formula
$\mathbb{A}_{t}=\frac{1}{16}\left(4s_{a}s_{b}-(s_{a}+s_{b}-s_{c})^{2}\right).$ (5)
For a tetrahedron $\tau\equiv\Delta(v_{0}v_{1}v_{2}v_{3})$ with squared edge lengths $\{s_{ij}\}_{0\leq i<j\leq 3}$, the squared volume is given by the Cayley–Menger determinant
$\mathbb{V}_{\tau}\;=\;\frac{1}{(3!)^{2}\,2^{3}}\begin{vmatrix}0&1&1&1&1\\ 1&0&s_{01}&s_{02}&s_{03}\\ 1&s_{01}&0&s_{12}&s_{13}\\ 1&s_{02}&s_{12}&0&s_{23}\\ 1&s_{03}&s_{13}&s_{23}&0\end{vmatrix}$. (6)
The squared area and volumes can alternatively be determined from the determinant of the corresponding Gram matrices [12, 61].
The complex dihedral angles in Eq. (3) capture the definitions of both Euclidean and Lorentzian angles. It naturally generalizes planar angles to higher dimensions. For a timelike edge within a Lorentzian tetrahedron, the corresponding dihedral angle is defined via Euclidean angles given by $\theta_{e,\tau}^{E}:=-\Theta_{e,\tau}\in[0,\pi]$. For a spacelike edge, the complex dihedral angle defines the corresponding Lorentzian dihedral angle which takes the form $\theta_{e,\tau}^{L}:=\Theta_{e,\tau}=\imath\beta-m\pi/2$, where $\beta\in\mathbb{R}$ and $m\in\{0,1,2\}$. The imaginary part $\beta$ corresponds to the rapidity between the adjacent face normals, characterizing the hyperbolic angle associated with the Lorentz boost. The integer $m$, which contributes to the real part of the dihedral angle, counts the number of lightrays that pass through the convex wedge formed by the two adjacent faces meeting at $e$. This wedge lies in the 2D plane orthogonal to the edge $e$ where the dihedral angle is defined. Geometrically, $m$ contributes to the causal structure of the edge by counting the number of lightrays intersecting its orthogonal plane. These lightrays originate from a point obtained by ‘projecting out’ the edge $e$, i.e., by collapsing all points along $e$ to a single vertex. In this projected two-dimensional geometry, the local lightrays and their intersections become manifest, allowing one to determine $m$ directly. Further details on the construction of this projection are given in Sections 3.3 and 3.4.
@@ 3 Lightcone analysis in Lorentzian triangulation
A defining feature of a Lorentzian geometry is the presence of a lightcone structure, which locally divides the tangent space at each spacetime point into causally distinct regions, i.e., timelike, spacelike and null (see Figure 1(a)). In the continuum, this structure is determined by the Lorentzian metric $g$ on a smooth manifold $(\mathcal{M},g)$: at any point $p\in\mathcal{M}$, the lightcone consists of all tangent vectors $X\in T_{p}\mathcal{M}$ satisfying $g(X,X)=0$. This causal decomposition underlies the notions of causal ordering, locality, and time orientation in spacetime.
@@
(b) A Lorentzian tetrahedron $\Delta(v_{0}v_{1}v_{2}v_{3})$ with its reference vertex $v_{0}$ at the origin intersecting the lightcones at $v_{0}$ (in orange). The intersection traces out conic sections on the opposite triangle $\Delta(v_{1}v_{2}v_{3})$.
Figure 1: Lightcones at a vertex in (2+1)–dimensional Minkowski spacetime.
3.1 Causal structure of a simplex
In a piecewise-flat Lorentzian triangulation, the metric is not given explicitly. Instead, the simplicial geometry is encoded in the set of edge lengths $s_{ij}$} assigned to the $1$-simplices. Each $d$-simplex is intrinsically flat and can therefore be embedded into a patch of Minkowski spacetime $\mathbb{R}^{1,d-1}$ with the standard metric $\eta=\mathrm{diag}(-1,1,\dots,1)$. The embedding, determined uniquely up to isometry, allows the reconstruction of the local causal structure of the simplex with respect to any of its vertices.
Definition 2 (Causal structure of a simplex relative to a vertex).
Let $\sigma^{d}\subset\mathbb{R}^{1,d-1}$ be a Lorentzian $d$-simplex with vertices $\{v_{0},v_{1},\dots,v_{d}\}$ and squared edge lengths $s_{ij}$}. The causal structure of $\sigma^{d}$ relative to the vertex $v_{0}$, denoted $\mathcal{C}(\sigma^{d};v_{0})$, is the partition of $\sigma^{d}$ induced by the two null hypersurfaces defined by
$\mathcal{N}^{\pm}(v_{0})\;=\;\{x\in\sigma^{d}\mid\eta_{\mu\nu}(x-v_{0})^{\mu}(x-v_{0})^{\nu}=0,\;\pm(x^{0}-v_{0}^{0})>0\}.$ (10)
The intersection of $\mathcal{N}^{\pm}(v_{0})$ with the interior and boundary of $\sigma^{d}$ determines the causal character (spacelike, timelike, or null) of regions within $\sigma^{d}$ relative to $v_{0}$. Two simplices $\sigma^{d}$ and $\tilde{\sigma}^{d}$ are said to have the same causal type at a vertex $v_{0}$ if their lightcone intersection patterns are topologically equivalent.
Thus, $\mathcal{C}(\sigma^{d};v_{0})$ encodes how the lightcones centered at $v_{0}$ intersect the simplex, distinguishing regions that are timelike, spacelike, or lightlike relative to the vertex (see Figure 1(b)). Throughout the paper, we shall refer to such regions simply as timelike, spacelike or lightlike, always understood to be with respect to the reference vertex. This provides a discrete analogue of the continuum causal structure induced by the local metric.
3.2 Causal structure in a Lorentzian triangulation
In a dd–dimensional Lorentzian triangulation, each vertex $v$ is shared by a collection of dd–simplices forming its star, denoted ${\rm St}(v)$. Topologically, ${\rm St}(v)$ is homeomorphic to a closed dd–ball, and its boundary $\partial{\rm St}(v)$ is a triangulated (d1)(d-1)–sphere composed of the faces opposite to $v$.
Definition 3 (Causal structure of a vertex neighbourhood).
Let $v$ be a vertex in a Lorentzian triangulation, and let ${\rm St}(v)$ denote the collection of dd–simplices $\sigma^{d}_{i}$ incident on $v$. The causal structure of the vertex neighbourhood of $v$, denoted $\mathcal{C}({\rm St}(v))$, is the collection
$\mathcal{C}({\rm St}(v))\;=\;\{\,\mathcal{C}(\sigma^{d}_{i};v)\ :\ \sigma^{d}_{i}\in{\rm St}(v)\,\},$
where each $\mathcal{C}(\sigma^{d}_{i};v)$ describes the intersection of the lightcones $\mathcal{N}_{i}^{\pm}(v)$ within the corresponding simplex $\sigma^{d}_{i}$.
Unlike in the continuum, where the local lightcones are directly defined by the metric tensor, piecewise-flat triangulations do not explicitly encode causality conditions. Instead, the causal relations must be reconstructed from the edge lengths. In this section, we present a combinatorial and algebraic framework for extracting the lightcone structure at a vertex in a (2+1)–dimensional Lorentzian triangulation directly from the squared edge lengths of the incident tetrahedra. We shall mainly restrict attention to Lorentzian simplices with non-degenerate subsimplices, that is, simplices without null edges or null faces, so that each edge and triangle has non-zero length or area. Null faces correspond to degenerate configurations where the induced metric becomes singular and their volumes vanish. Their inclusion requires a separate treatment of signature transitions and degenerate Gram matrices, considerably complicating the classification. Moreover, in the Regge calculus, null hinges do not contribute directly to the action since the associated volumes vanish. For these reasons, we treat simplices with null faces as limiting cases separating distinct causal types and leave a systematic treatment of such degenerate configurations to future work.
3.3 Local vertex–causality conditions in a triangulation